Bartello inc. provided the following information related to copyright it owns at December 31, 2017. Bartello will continue to use the copyright in the future.

Cost $8,600,000

Carrying amount 4,300,000

Recovery amount 3, 400, 000

Required;

  1. Prepare the journal entry (if any) to record the impairment of the assets at December 31, 2017.
  2. Prepare the journal entry to record Amortization expense for 2017 related to copyright.
  3. The fair value of the copyright at December 312, 2017is $3,500,000. Prepare the journal entry (if any) necessary to record the change in fair value.
Answer #1

Impairment of asset is recorded when carrying amount of asset exceed recoverable amount.

Impairment = carrying value - recoverable amount = 43,00,000 - 34,00,000 = 9,00,000

journal entry

date entry debit credit
31 dec impairment loss 900000
copyright 900000
(impairment loss recorded)

Ammortization expense

date entry debit credit
31 dec ammotization 4300000
copyright 4300000

(8600000-4300000= 4300000)

ammortization on copyright recorded

No entry shall be made for fair value of copyright as recoverable amount is higher of fair value minus cost to sale the asset and the asset is now recorded on recoverable value so this entry is not required